# This work has been supported by the Ministerio de Ciencia, Innovación y Universidades of the Spanish Government under HVD (PID2021-125051OB-I00) and SEGA-CV (TED2021-131643A-I00) projects, and the Comunidad Autónoma de Madrid under the Grant IND2020/TIC-17515

### v2

@@ 4.1 Pose Estimation Experimental Settings
In all our experiments we consistently use the same architecture and the same evaluation metrics, described in Section 4.1.1 and Section 4.1.2 respectively. All the models described are trained using PyTorch [22] with input 512x512 grayscale images. We employ the Adam optimiser [23] with a learning rate of $0.0001$. The ground-truth heatmaps are created with SPIN always using the same parameters (sigma deviation of 7 pixels) and all training parameters are kept the same for all models, including the random seed. This approach is adopted to reduce the influence of factors other than the input training data.
In the SPEED+ experiments, we train two distinct models: the first on the original synthetic SPEED+ training split, which comprises 47,966 images. The second on the corresponding images replicated using the SPIN tool. Both models are trained for 60 epochs and are tested over the testbed domains of SPEED+: Sunlamp with 2,791 images, and Lightbox with 6,740.
For the SHIRT experiments, we train four separate models. Two are trained on the original synthetic sequences ROE1 and ROE2, each containing 2,371 images. The other two models are trained on replicas of these sequences, created using the SPIN tool. These four models are trained for 5 epochs and are evaluated in the testbed domains of ROE1 and ROE2. These testbed domains consist of the same poses as those in their respective training sequences, but the testing is conducted in a different environment specific to the testbed domains.
4.1.1 Pose Estimation Model
@@ 4.1.2 Metrics
Table 3: Comparison of spacecraft pose estimation performance between the original SPEED+ dataset and the SPEED+ replica generated with SPIN. Lower scores indicate better performance.
Training data Test $S_{v}$ $S_{q}$ $S$
SPEED+ Lightbox 0.359 1.361 1.715
Sunlamp 0.356 1.665 2.021
SPIN Lightbox 0.171 0.719 0.891
Sunlamp 0.122 0.621 0.743
Table 4: Comparison of spacecraft pose estimation performance between the original SHIRT dataset over the ROE1 and ROE2 sequences and the respective replicas generated with SPIN.
ROE $S_{v}$ $S_{q}$ $S$
SHIRT 1 0.072 0.542 0.614
2 0.526 2.131 2.021
SPIN 1 0.030 0.352 0.381
2 0.176 0.987 1.163
4.2 SPIN Generation Settings
@@
For the generation of the dataset replicas with SPIN we set some parameters to match those of SPEED+ and SHIRT with the aim of keeping a consistent input domain, and we set others to improve the image realism. We provide a summary of the modified parameters in Table 2. Specifically, for camera settings, we keep the intrinsics, color adjustment, and noise parameters constant. This ensures that the spacecraft is viewed from the same perspective and that noise levels are uniform, for fair comparison. However, we modify glare and bloom settings to enhance the scene’s realism. For the environment settings, we use shadow rendering techniques similar to those used in the SPEED+ dataset. To minimize possible effects of Earth’s presence in the background for the task of pose estimation, we choose to use similar background images in our synthetic dataset as those in SPEED+. Regarding illumination, we adjust the settings to produce more realistic images, in particular, we employ the Sunlight setting. Lastly, we activate the high-quality materials to create images that more closely resemble those from the testbed domain. In Figure 7 we display examples of the differences between images generated using a configuration aligned with the SPEED+ synthetic subset, and the same images after applying our enhancement settings.
Figure 5: Examples of the estimated pose represented as a wireframe model (solid blue for SPIN and dashed yellow for SPEED+) overlaid over the real satellite for SPEED+ images of the Lightbox (left) and Sunlamp (right) testbeds.
4.3 Quantitative Results
Table 3 indicates that algorithms trained on our synthetic dataset exhibit significant improvements over those trained on the SPEED+ dataset: a 48% reduction in the error rates for the Lightbox testbed, moving from a baseline error of 1.71 to a 0.891; and a 63% drop for the Sunlamp one, from 2.021 to 0.743. For the SHIRT dataset, results in Table 4 indicate a 47% error rate reduction for ROE1, and a 53% one for ROE2. Visual examples of the pose estimation errors are presented in Figure 5.
These experiments show how SPIN allows to reduce the performance gap between the synthetic and testbed domains, by providing more realistic training images than those of the existing synthetic datasets from the literature.
4.3.1 Ablation Tests
In this section, we include an ablation study to evaluate the specific impact of the Camera, Environment, and Material settings modifications for improved realism on the pose estimation performance. We conduct the study by replicating different synthetic versions of the training split of the SPEED+ with SPIN. Each replica is created with a different combination of SPIN settings and used to train a separate pose estimation model. We illustrate the effect of the SPIN settings on a sample image in Figure 6. As in previous experiments, we evaluate over the testbed SPEED+ domains of Sunlamp and Lightbox.
Figure 6: Comparison of a synthetic SPEED+ image and those generated with SPIN under different configurations: Baseline indicates that no setting is activated, and ”C,” ”E,” and ”M” indicate that settings for the Camera, Environment, and Material are enabled, respectively.
Table 5: Impact of SPIN settings (indicated by a checkmark ($\checkmark$) when activated and a dash (-) when deactivated) on the pose estimation performance over the SPEED+ testbeds. We activate or deactivate only those parameters that are configured differently from SPEED+, as indicated in Table 2.
@@ 4.3.1 Ablation Tests
Camera Environment Materials $S_{v}$ $S_{q}$ $S$ $S_{v}$ $S_{q}$ $S$
- - - 0.291 1.260 1.551 0.379 1.808 2.187
$\checkmark$ - - 0.352 1.450 1.803 0.416 1.793 2.209
- $\checkmark$ - 0.309 1.269 1.578 0.224 1.146 1.370
- - $\checkmark$ 0.381 1.483 1.864 0.437 1.775 2.213
$\checkmark$ $\checkmark$ - 0.217 0.911 1.13 0.132 0.759 0.891
- $\checkmark$ $\checkmark$ 0.353 1.508 1.861 0.392 1.657 2.049
$\checkmark$ - $\checkmark$ 0.270 1.093 1.364 0.215 1.103 1.319
$\checkmark$ $\checkmark$ $\checkmark$ 0.171 0.719 0.891 0.122 0.621 0.743
Table 5 summarises the results. We first define a baseline –first row– by deactivating all settings. Next, we evaluate each setting independently. Activating each setting independently does not result in significant performance improvements, probably due to the lack of an intense directional light to trigger camera effects such as lens bloom or better reflective materials, hence generating images very similar to the baseline ones. The activation of the Environment settings leads to the most effective results, allowing for the generation of more challenging images, as the shadows cast by the directional light occlude parts of the satellite.
Combining couples of settings yields better results. Interestingly, when Materials and Environment settings are activated, while results improve with respect to just activating Materials, they do not exceed those of just activating Environment. Best results are clearly obtained by combining Camera and Environment. This synergy might be attributed to the increased illumination on the spacecraft due to either increased reflectivity or to a more intense light source, which coupled with camera settings such as bloom, results in much more realistic images. It is noteworthy, however, that this same reflectivity and intense illumination can diminish performance when these camera settings are not enabled, leading to overexposed images.
Finally, enabling all three settings results in the best performance, reaching the results presented in Section 4.3.
5 Conclusions
In conclusion, SPIN significantly enhances realism in image generation, narrowing the gap between synthetic and real imagery in pose estimation compared to existing synthetic datasets. It also provides a wider range of ground-truth data, including the use of dense pose labels, that no other dataset in the literature provides. Moreover, SPIN facilitates the creation of new and diverse test scenarios, expanding the variety and depth of ground-truth data in current datasets.
Figure 7: Image comparison between a baseline configured like the synthetic split from the SPEED+ dataset, and the same images after applying Environment, Camera, and Material settings.

### v3

@@ IV-A Pose Estimation Experimental Settings
In all our experiments we consistently use the same architecture and the same evaluation metrics, described in Section IV-A1 and Section IV-A2 respectively. All the models described are trained using PyTorch [21] with input 512x512 grayscale images. We employ the Adam optimiser [22] with a learning rate of $0.0001$. The ground-truth heatmaps are created with SPIN always using the same parameters (sigma deviation of 7 pixels) and all training parameters are kept the same for all models, including the random seed. This approach is adopted to reduce the influence of factors other than the input training data.
IV-A1 Pose Estimation Model
@@ IV-A2 Metrics
IV-B Generation Settings
@@
To generate our version of the SPEED+[1] dataset using SPIN, we set some parameters to match those of SPEED+ and SHIRT[9] with the aim of keeping a consistent input domain. Specifically, for camera settings, we keep the intrinsics, color adjustment, and noise parameters constant. This ensures that the spacecraft is viewed from the same perspective and that noise levels are uniform, for fair comparison. However, we modify glare and bloom settings to enhance the scene’s realism. For the environment settings, we use shadow rendering techniques similar to those used in the SPEED+ dataset. To minimize possible effects of Earth’s presence in the background for the task of pose estimation, we choose to use similar background images in our synthetic dataset as those in SPEED+. Regarding illumination, we adjust the settings to produce images with more intense directional lighting and harder shadows, resembling those that can be found on the testbed domains.
IV-C Quantitative Results
To validate our dataset, we have trained the model from [23] with our SPIN-generated dataset and with SPEED+, with the latter being used as a baseline. For both datasets, the same hyper-parameters were used, training the model for 60 epochs (120000 iterations), using a batch size of 24 images scaled to 512x512, and with a learning rate of $2.5\times 10^{-4}$.
In Table II we summarize the performance results of a model trained on SPEED+ Synthetic train subset and a model trained on our SPIN-generated dataset with the same train poses. Models are evaluated on both testbed subsets from SPEED+ (Lightbox and Sunlamp), and also on the testbed subsets from SHIRT (ROE1 and ROE2). We can see that the dataset generated with SPIN is able to replicate the dataset successfully. Not only that, but our more challenging training examples result in a better generalization across testbed images from both SPEED+ and SHIRT, with a 45% reduction in the error rates for the Lightbox testbed, reducing the error from 1.764 to a 0.962, and a 64% for the Sunlamp subset, from 2,083 to 0.760; along an error reduction of 36% from 2.786 to 1.788 for the ROE1 testbed subset, and a 42% for the ROE2 subset, from 2.450 to 1.421, validating our simulator dataset generation capabilities.
IV-D Exploiting SPIN: In-scene data augmentation
SPIN’s primary advantage over conventional datasets lies in its versatility. It enables users to generate unlimited data, opening the door to new methods and techniques that are not feasible with traditional datasets. In computer vision, data augmentation techniques are commonly employed to significantly alter image appearances, creating diverse, unseen data to enhance deep learning model training. Our simulator achieves this goal more effectively. Instead of augmenting existing 2D images, we perform scene augmentation. This approach ensures pixel-perfect preservation of image structure while dramatically changing its appearance. For instance, we can modify illumination intensity and origin, spacecraft materials, or background. We call our proposed method in-scene data augmentation.In Fig. 4 we show an example of the same pose augmented with our simulation tool using different rendering and lightning settings.
Figure 4: In-Scene Data Augmentation example. Different renderings for the same pose generated with our tool with different combinations of background, illumination, and color settings.
We evaluate the effectiveness of our method by training using three different synthetic appearances for each pose in the training dataset. We use the same hyper-parameters from our validation test, except we now train for 20 epochs to ensure that we train for the same amount of iterations (120000). Results can be found in Table II. We can see how SPIN capabilities to generate in-scene data augmentation help the model generalize better in both SPEED+ testbed subsets, with an average 21% error reduction, and also in SHIRT testbed subsets, with a 27% error reduction.
TABLE II: Spacecraft pose estimation performance after training on the SPEED+ dataset and our SPIN-generated SPEED+ replicas. ISDA stands for in-scene data augmentation.
Test Set Training Set $S_{v}\left(\downarrow\right)$ $S_{q}\left(\downarrow\right)$ $S\left(\downarrow\right)$
Lightbox SPEED+ 0.363 1.401 1.764
SPIN 0.179 0.783 0.962
ISDA 0.143 0.614 0.757
Sunlamp SPEED+ 0.382 1.701 2.083
SPIN 0.118 0.642 0.760
ISDA 0.104 0.490 0.594
ROE1 SPEED+ 0.665 2.121 2.786
SPIN 0.225 1.563 1.788
ISDA 0.155 1.202 1.357
ROE2 SPEED+ 0.502 1.948 2.450
SPIN 0.283 1.138 1.421
ISDA 0.200 0.797 0.996
IV-E Ablation Test
We conducted an ablation study to measure the impact of different SPIN settings on dataset generation and model performance. This study was motivated by our observation that the algorithm trained on our SPIN-generated replica of SPEED+ generalized better to testbed domains. We created several versions of SPEED+ using various combinations of SPIN settings. These settings were categorized into three groups: Camera settings (Glare/Bloom enabled), Environment settings (light intensity and shadows enabled), and Material settings (high-quality, reflective materials). We then measured how each setting affected the model’s performance.
Table III presents the results of the ablation study. The baseline—shown in the first row—represents the configuration with all settings disabled. We then evaluate the effect of enabling each setting individually. Activating camera effects significantly enhances performance, especially within the Sunlamp subset. While enabling environment settings also boosts performance, this improvement is observed only in the Sunlamp domain. Interestingly, activating materials alone leads to a performance drop in the Lightbox subset and offers no advantage in Sunlamp. Although the environment and material settings do not independently improve performance, their combination with camera effects shows a marked improvement, demonstrating the cumulative benefits of these settings when used together, ultimately enhancing the model trained on these dataset versions.
TABLE III: Impact of SPIN settings pose estimation performance over the SPEED+ testbeds. The C,E,M letters stand for Camera, Environment and Material respectively.
@@ IV-E Ablation Test
C E M $S_{v}\left(\downarrow\right)$ $S_{q}\left(\downarrow\right)$ $S\left(\downarrow\right)$ $S_{v}\left(\downarrow\right)$ $S_{q}\left(\downarrow\right)$ $S\left(\downarrow\right)$
- - - 0.305 1.295 1.600 0.398 1.886 2.254
$\checkmark$ - - 0.289 1.177 1.466 0.188 1.045 1.233
- $\checkmark$ - 0.312 1.293 1.604 0.244 1.86 1.430
- - $\checkmark$ 0.395 1.505 1.899 0.468 1.767 2.236
$\checkmark$ $\checkmark$ - 0.224 0.976 1.200 0.139 0.760 0.899
- $\checkmark$ $\checkmark$ 0.344 1.527 1.871 0.394 1.711 2.105
$\checkmark$ - $\checkmark$ 0.254 1.060 1.314 0.204 1.040 1.244
$\checkmark$ $\checkmark$ $\checkmark$ 0.179 0.783 0.962 0.118 0.642 0.760
V Conclusions
In conclusion, SPIN offers a new tool to generate datasets with improved visuals, narrowing the gap between synthetic and real imagery in pose estimation compared to existing synthetic datasets. It also provides a wider range of ground-truth data, including the use of dense pose labels, that no other dataset in the literature provides. Moreover, SPIN facilitates the creation of new and diverse test scenarios, expanding the variety and depth of ground-truth data in current datasets.